# Beyond Ten Turns: Unlocking Long-Horizon Agentic Search with Large-Scale Asynchronous RL

### v3

@@ Abstract
Recent advancements in LLM-based agents have demonstrated remarkable capabilities in handling complex, knowledge-intensive tasks by integrating external tools. Among diverse choices of tools, search tools play a pivotal role in accessing vast external knowledge. However, open-source agents still fall short of achieving expert-level Search Intelligence, the ability to resolve ambiguous queries, generate precise searches, analyze results, and conduct thorough exploration. Existing approaches fall short in scalability, efficiency, and data quality. For example, small turn limits in existing online RL methods, e.g. $\leq 10$, restrict complex strategy learning. This paper introduces ASearcher, an open-source project for large-scale RL training of search agents. Our key contributions include: (1) Scalable fully asynchronous RL training that enables long-horizon search while maintaining high training efficiency. (2) A prompt-based LLM agent that autonomously synthesizes high-quality and challenging QAs, creating a large-scale QA dataset. Through RL training, our prompt-based QwQ-32B agent achieves substantial improvements, with 46.7% and 20.8% Avg@4 gains on xBench and GAIA, respectively. Notably, our agent exhibits extreme long-horizon search, with tool calls exceeding 40 turns and output tokens exceeding 150k during training time. With a simple agent design and no external LLMs, ASearcher-Web-QwQ achieves Avg@4 scores of 42.1 on xBench and 52.8 on GAIA, surpassing existing open-source 32B agents.  [^1] We open-source our models, training data, and codes in https://github.com/inclusionAI/ASearcher.
Figure 1: (Left) Asynchronous RL brings substantial improvements: Through RL training, our agent, ASearcher-Web-QwQ, obtains 20.8%, 46.7%, and 20.4% improvements on GAIA, xBench, and Frames, respectively. (Middle) & (Right) Through RL training, ASearcher-Web-QwQ learns to conduct long-horizon search, with tool calls exceeding 40 turns and output tokens exceeding 150k during training. The agent also learns expert-level search strategies (See case study in Sec. 2)
@@ 1 Introduction
Recent advances in LLM-based agents have demonstrated remarkable capabilities in solving complex, knowledge-intensive problems by leveraging single or multiple external tools Xi et al. (2025); Yao et al. (2023); Wang et al. (2024a). Among these, search tools stand out as particularly critical, enabling agents to access vast external knowledge for enhanced problem-solving OpenAI (2025); Google Team (2025); Perplexity Team (2025). However, expert-level use of search requires advanced intelligence. For instance, consider the question “As of December 31, 2024, what were the numbers of gold, silver, and bronze medals won by China in the 2012 London Olympics?”.While seemingly straightforward, this query is indeed challenging due to conflicting answers online (e.g., “38 gold, 27 silver, 22 bronze” vs. “39 gold, 31 silver, 22 bronze”). A search agent must navigate noisy and conflicting answers from diverse sources, identify the root cause of conflicts as doping test disqualifications from official reports, and ultimately determine the correct answer. Challenging real-world tasks require the agent to resolve high uncertainty in input queries, generate precise search queries, analyze and extract key insights from massive data, resolve inconsistencies, and conduct in-depth exploration. We term this advanced capability "Search Intelligence".
Proprietary agents and models has already exhibit signs of complex search behaviors through large-scale Reinforcement Learning (RL) training (AI, 2025; OpenAI, ). However, open-source approaches for developing search agents still face significant limitations. A series of works employ Reinforcement Learning or Supervised Fine-Tuning approaches to incentivize tool-using capabilities (Jin et al., 2025; Song et al., 2025; Zheng et al., 2025; Tan et al., 2025). On the other hand, prompt-based LLM agents supported by open-source models could perform massive tool calls without training (Li et al., 2025c; Alzubi et al., 2025). However, in practice, we find that existing online RL approaches fail to incentivize complex and effective search strategies. We also find prompt-based LLM agents could fail due to the insufficient capabilities of the LLM, such as failing to precisely extract key information from noisy webpages and unable to verify wrong conclusions. More recently, some works further build up on prompt-based LLM agents, utilizing offline RL approaches to improve the prompt-based agents (Sun* et al., 2025; Li et al., 2025d). However, this offline RL paradigm, has been shown to underperform online RL in a broader range of domains (Xu et al., 2024; Fu et al., 2021; Song et al., 2024).
In reasoning tasks such as math and coding, online RL has enable the models to evolve complex behaviors through iterative refining the reasoning processes based on correctness feedback. Guo et al. (2025); Luo et al. (2025b); Fu et al. (2025),. This raises a critical question: How could online RL methods effectively unlock Search Intelligence in open-source agents?
@@
  • Insufficient search turns limit complex strategy learning. Existing works, such as Search-R1 Jin et al. (2025), artificially limit the number of search turns, e.g. $\leq 10$ per trajectory, preventing the agent from exploring deeper search paths. However, complex queries often require multi-turn tool calls and multi-step reasoning, that could not be learned under strict turn limits.
  • Lack of large-scale, high-quality question-answer (QA) pairs: RL training for reasoning tasks requires abundant, challenging, and correct QA pairs An et al. ; Li et al. (2025a); Yu et al. (2025). However, most existing open-source datasets for search agents are often outdated (e.g. HotpotQA), oversimplified, or too small, failing to stimulate complex search behaviors through RL Yang et al. (2018); Li et al. (2025b); Tao et al. (2025).
@@
  • Long-horizon search via fully asynchronous agentic RL training. With a large turn limit in batch generation RL training systems (Jin et al., 2025; Song et al., 2025; Luo et al., 2025a; Team et al., 2025), long trajectories within a batch could easily lead to significant idle time, slowing down the whole training process. Building up on AReaL  (Fu et al., 2025), our fully asynchronous system avoids long trajectories from blocking the training by decoupling trajectory execution from model updates. This allows relaxed turn limits (e.g., 128 turns/trajectory), enabling agents to explore deeper search paths without sacrificing training efficiency. Remarkably, our agent, ASearcher-Web-QwQ, achieves extreme long-horizon search, with tool calls exceeding 40 turns and generated tokens surpassing 150k during RL training.
@@
We evaluate our agents with on multi-hop QA benchmarks and challenging benchmarks including GAIA (Mialon et al., 2023) , xbench-DeepSearch (Xbench-Team, 2025), and Frames (Krishna et al., 2024). ASearcher-Local-7B/14B, trained only with local knowledge base, demonstrate surprisingly generalizability to realistic web search and achieve state-of-the art performances on multi-hop and single-hop QA tasks. Building up on QwQ-32B, ASearcher-Web-QwQ achieves an Avg@4 score of 42.1 on xBench-DeepSearch and 52.8 on GAIA, surpassing a set of open-source agents. When evaluating Pass@4, ASearcher-Web-QwQ achieves 70.1 on GAIA and 68.0 on xBench-DeepSearch. Notably, through RL training, ASearcher-Web-QwQ obtains 46.7% and 20.8% improvements on xBench-DeepSearch and GAIA, respectively.
@@ 2 Limitations of Existing Open-source Approaches
In this section, we provide a detailed case study on an extremely challenging question from GAIA (Mialon et al., 2023). Specifically, we analyze Search-R1-32B (Jin et al., 2025) and Search-o1 (QwQ) (Li et al., 2025c) in Fig. 3. The detailed trajectories are provided in Appendix A.
@@ 3.1 Agent Design
  • For base LLMs, we following prior works (Jin et al., 2025; Song et al., 2025), to adopt append-only style prompting for the agent. Specifically, starting from a system prompt, all LLM-generated responses, search results and summaries of webpages are appended to the history. The agent takes as input the full history in chronological order and outputs some reasoning texts and actions. This approach ensures efficiency during inference time.
@@ 3.2 Training Data
We begin with the training sets from HotpotQAYang et al. (2018) and 2WikiMultiHopQAHo et al. (2020), both of which are multi-hop QA datasets. We employ a model-based filtering process. We first train a model on the full set of open-source data with RL, and then generate 16 responses for each question using the trained model. Finally, we filter out questions that meat any of the following criteria,
@@
Additionally, we include a set of question-answer (QA) pairs designed for accessing certain webpages. In particular, we incorporate a small subset of WebWalkerQAWu et al. (2025b) to help the model learn how to locate answers within noisy, real-world web search environments.
@@ 3.3.1 Challenges of Scaling Up Trajectory Length in RL
Agentic tasks often require extensive LLM generations and multiple tool calls to solve complex problems, leading to prolonged trajectory execution time. As shown in Fig. 6(Left), we evaluate our RL-trained QwQ-32B agent on GAIA Mialon et al. (2023), xBench-Deepsearch Krishna et al. (2024) and Frames (Krishna et al., 2024), forcing the agent to use tools for different minimal turn numbers. The results demonstrate that accuracy improves with more turns, confirming that complex tasks demand longer trajectories for effective problem-solving.
@@
Both prolonged execution and high runtime variance degrade RL training efficiency. We take one-step-off RL training system (Luo et al., 2025a) as a representative example for batch generation RL systems. In one-step-off RL training, training for step N and trajectory generation for step N+1 are executed concurrently. As shown in Fig. 7, though this system overlaps trajectory rollouts with model training, batch generation remains bottlenecked by the slowest trajectory (e.g., trajectory 7), causing GPU idle time and under-utilization.
@@ 3.4 Training Details
We employ the GRPO (Shao et al., 2024) algorithm to train search agents. Specifically, for each input question $x$, $G$ trajectories $\tau_{1},\tau_{2},\cdots,\tau_{G}$ are generated where $\tau_{i}=(s_{0}^{i},a^{i}_{0},s_{1}^{i},\cdots,s_{T_{i}}^{i})$. To optimize the agent, we employ the following loss,
@@
For reward function, we adopt a sparse-reward setting where rewards are computed at trajectory completion. When training from base LLMs, the reward function combines a format reward and F1 score through multiplication. When fine-tuning LRM-based agents (e.g., QwQ), we utilize LLM-as-JudgeLiu et al. (2023)Wang et al. (2024b) as the reward function and omit format rewards, as these models inherently maintain proper output formatting.
@@ 4.1 Experiment Setup
We first evaluate the agents on single-hop and multi-hop QA tasks. For single-hop questions, we use Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017) and PopQA (Mallen et al., 2022). For multi-hop questions, we use HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (Ho et al., 2020), MuSiQue (Trivedi et al., 2022), and Bamboogle (Press et al., 2022). We further perform evaluation on more challenging benchmarks including Frames (Krishna et al., 2024), GAIA (Mialon et al., 2023), and xBench-DeepSearch (Xbench-Team, 2025) as extra test sets. We evaluate our approach on 1000 randomly sampled instances from the validation sets of HotpotQA, 2WikiMultiHopQA, and MuSiQue. For Bamboogle, Frames, GAIA and xBench-DeepSearch, we use their full test sets. For GAIA, we use the 103 examples from the text-only validation subset (Li et al., 2025c).
@@
We evaluate the search agents with two settings, each with different types of search tools. In the first setting, local knowledge base with RAG, agents interact with a locally deployed RAG system to retrieve related information from a Wikipedia 2018 corpus (Karpukhin et al., 2020). In the other web-based search and browsing setting, agents operate in an interactive web environment with access to both a search engine and a browser tool. For more challenging benchmarks, GAIA, xBench-DeepSearch and Frames, we only conduct evaluations under this web-based setting.
@@
We consider two groups of baselines aligned with the two benchmark categories. For the multi-hop and single-hop QA benchmarks, we include Search-R1(7B/14B/32B) (Jin et al., 2025), R1-Searcher(7B) (Song et al., 2025), Search-o1(QwQ-32B) (Li et al., 2025c), DeepResearcher (Zheng et al., 2025) and SimpleDeepSearcher (Sun* et al., 2025). We also prompt Qwen-2.5-7B/32B to directly generate answers without using any tools. On the more challenging benchmarks, we compare against powerful 32B-scale models, including direct generation with QwQ-32B, Search-o1(QwQ-32B) (Li et al., 2025c), Search-R1-32B (Jin et al., 2025), WebThinker-QwQ (Li et al., 2025d),SimpleDeepSearcher-QwQ (Sun* et al., 2025) and WebDancer-32B (Wu et al., 2025a). All baselines are evaluated using the same tools as our agent to ensure a fair comparison.
@@
We set the turn limit as 32 for 7B and 14B models, and 128 for ASearcher-Web-QwQ. The batch size is set as 128 for 7B and 14B models, and 64 for ASearcher-Web-QwQ. We curate two sets of training data, one for 7B/14B training and the other for QwQ-32B training. These two datasets are both of 35k sizes and open-sourced. Training of ASearcher-Web-QwQ takes approximated 7.6k H800 GPU hours.
@@ 4.2 Main Results
As shown in Fig. 8, ASearcher-Web-QwQ obtains +9.1, +13.4, and +12.0 improvements on GAIA, xBench-DeepSearch and Frames respectively. When considering the pass rate, i.e. Pass@4, ASearcher-Web-QwQ also obtains significant gains, especially on xBench-DeepSearch with 17.0 improvements. Significant improvements in pass rate demonstrate that our training pipeline trains the agent to learn complex search strategies to perform precise searches, extract key information, and resolve conflict information.

### v4

@@ Abstract
Recent advancements in LLM-based agents have demonstrated remarkable capabilities in handling complex, knowledge-intensive tasks by integrating external tools. Among diverse choices of tools, search tools play a pivotal role in accessing vast external knowledge. However, open-source agents still fall short of achieving expert-level Search Intelligence, the ability to resolve ambiguous queries, generate precise searches, analyze results, and conduct thorough exploration. Existing approaches fall short in scalability, efficiency, and data quality. For example, small turn limits in existing online RL methods, e.g. $\leq 10$, restrict complex strategy learning. This paper introduces ASearcher, an open-source project for large-scale RL training of search agents. Our key contributions include: (1) Scalable fully asynchronous RL training that enables long-horizon search while maintaining high training efficiency. (2) A prompt-based LLM agent that autonomously synthesizes high-quality and challenging QAs. Through RL training, our prompt-based QwQ-32B agent achieves substantial improvements. Notably, our agent exhibits extreme long-horizon search, with tool calls exceeding 100 turns and output tokens exceeding 400k during training time. With a simple agent design and no external LLMs, ASearcher-Web-QwQ achieves Avg@4 scores of 51.1 on xBench and 58.7 on GAIA, surpassing existing open-source 32B agents. Finally, we also show that ASearcher-Web-QwQ could achieve performance of commercial systems using external summary tool in a zero-shot transfer manner and test-time search.  [^1] We open-source our models, training data, and codes in https://github.com/inclusionAI/ASearcher.
Figure 1: (Left) Asynchronous RL brings substantial improvements: Through RL training, our agent, ASearcher-Web-QwQ, obtains +15.0, +22.4, and +15.6 improvements on GAIA, xBench, and Frames, respectively. (Middle) & (Right) Through RL training, ASearcher-Web-QwQ learns to conduct long-horizon search, with tool calls exceeding 100 turns and output tokens exceeding 400k during training. The agent also learns expert-level search strategies (See case study in Sec. 2)
@@ 1 Introduction
Recent advances in LLM-based agents have demonstrated remarkable capabilities in solving complex, knowledge-intensive problems by leveraging single or multiple external tools [42, 45, 37]. Among these, search tools stand out as particularly critical, enabling agents to access vast external knowledge for enhanced problem-solving [26, 8, 27]. However, expert-level use of search requires advanced intelligence. For instance, consider the question “As of December 31, 2024, what were the numbers of gold, silver, and bronze medals won by China in the 2012 London Olympics?”.While seemingly straightforward, this query is indeed challenging due to conflicting answers online (e.g., “38 gold, 27 silver, 22 bronze” vs. “39 gold, 31 silver, 22 bronze”). A search agent must navigate noisy and conflicting answers from diverse sources, identify the root cause of conflicts as doping test disqualifications from official reports, and ultimately determine the correct answer. Challenging real-world tasks require the agent to resolve high uncertainty in input queries, generate precise search queries, analyze and extract key insights from massive data, resolve inconsistencies, and conduct in-depth exploration. We term this advanced capability "Search Intelligence".
Proprietary agents and models has already exhibit signs of complex search behaviors through large-scale Reinforcement Learning (RL) training [1, 25]. However, open-source approaches for developing search agents still face significant limitations. A series of works employ Reinforcement Learning or Supervised Fine-Tuning approaches to incentivize tool-using capabilities [11, 30, 49, 33]. On the other hand, prompt-based LLM agents supported by open-source models could perform massive tool calls without training [18, 2]. However, in practice, we find that existing online RL approaches fail to incentivize complex and effective search strategies. We also find prompt-based LLM agents could fail due to the insufficient capabilities of the LLM, such as failing to precisely extract key information from noisy webpages and unable to verify wrong conclusions. More recently, some works further build up on prompt-based LLM agents, utilizing offline RL approaches to improve the prompt-based agents [32, 19]. However, this offline RL paradigm, has been shown to underperform online RL in a broader range of domains [43, 6, 31].
In reasoning tasks such as math and coding, online RL has enable the models to evolve complex behaviors through iterative refining the reasoning processes based on correctness feedback. [9, 22, 7],. This raises a critical question: How could online RL methods effectively unlock Search Intelligence in open-source agents?
@@
  • Insufficient search turns limit complex strategy learning. Existing works, such as Search-R1 [11], artificially limit the number of search turns, e.g. $\leq 10$ per trajectory, preventing the agent from exploring deeper search paths. However, complex queries often require multi-turn tool calls and multi-step reasoning, that could not be learned under strict turn limits.
  • Lack of large-scale, high-quality question-answer (QA) pairs: RL training for reasoning tasks requires abundant, challenging, and correct QA pairs [3, 16, 46]. However, most existing open-source datasets for search agents are often outdated (e.g. HotpotQA), oversimplified, or too small, failing to stimulate complex search behaviors through RL [44, 17, 34].
@@
  • Long-horizon search via fully asynchronous agentic RL training. With a large turn limit in batch generation RL training systems [11, 30, 21, 35], long trajectories within a batch could easily lead to significant idle time, slowing down the whole training process. Building up on AReaL  [7], our fully asynchronous system avoids long trajectories from blocking the training by decoupling trajectory execution from model updates. This allows relaxed turn limits (e.g., 128 turns/trajectory), enabling agents to explore deeper search paths without sacrificing training efficiency. Remarkably, our agent, ASearcher-Web-QwQ, achieves extreme long-horizon search, with tool calls exceeding 100 turns and generated tokens surpassing 400k during RL training.
@@
We evaluate our agents with on multi-hop QA benchmarks and challenging benchmarks including GAIA [24] , xbench-DeepSearch [41], and Frames [14]. ASearcher-Local-7B/14B, trained only with local knowledge base, demonstrate surprisingly generalizability to realistic web search and achieve state-of-the art performances on multi-hop and single-hop QA tasks. Building up on QwQ-32B, ASearcher-Web-QwQ achieves an Avg@4 score of 51.1 on xBench-DeepSearch and 58.7 on GAIA, surpassing a set of open-source agents. When evaluating Pass@4, ASearcher-Web-QwQ achieves 74.7 on GAIA and 75.0 on xBench-DeepSearch. Notably, through RL training, ASearcher-Web-QwQ obtains 78.0% and 34.3% improvements on xBench-DeepSearch and GAIA, respectively.
@@ 2 Limitations of Existing Open-source Approaches
In this section, we provide a detailed case study on an extremely challenging question from GAIA [24]. Specifically, we analyze Search-R1-32B [11] and Search-o1 (QwQ) [18] in Fig. 3. The detailed trajectories are provided in Appendix A.
@@ 3.1 Agent Design
  • For base LLMs, we following prior works [11, 30], to adopt append-only style prompting for the agent. Specifically, starting from a system prompt, all LLM-generated responses, search results and summaries of webpages are appended to the history. The agent takes as input the full history in chronological order and outputs some reasoning texts and actions. This approach ensures efficiency during inference time.
@@ 3.2 Training Data
We begin with the training sets from HotpotQA[44] and 2WikiMultiHopQA[10], both of which are multi-hop QA datasets. We employ a model-based filtering process. We first train a model on the full set of open-source data with RL, and then generate 16 responses for each question using the trained model. Finally, we filter out questions that meat any of the following criteria,
@@
Additionally, we include a set of question-answer (QA) pairs designed for accessing certain webpages. In particular, we incorporate a small subset of WebWalkerQA[40] to help the model learn how to locate answers within noisy, real-world web search environments.
@@ 3.3.1 Challenges of Scaling Up Trajectory Length in RL
Agentic tasks often require extensive LLM generations and multiple tool calls to solve complex problems, leading to prolonged trajectory execution time. As shown in Fig. 6(Left), we evaluate our RL-trained QwQ-32B agent on GAIA [24], xBench-Deepsearch [14] and Frames [14], forcing the agent to use tools for different minimal turn numbers. The results demonstrate that accuracy improves with more turns, confirming that complex tasks demand longer trajectories for effective problem-solving.
@@
Both prolonged execution and high runtime variance degrade RL training efficiency. We take one-step-off RL training system [21] as a representative example for batch generation RL systems. In one-step-off RL training, training for step N and trajectory generation for step N+1 are executed concurrently. As shown in Fig. 7, though this system overlaps trajectory rollouts with model training, batch generation remains bottlenecked by the slowest trajectory (e.g., trajectory 7), causing GPU idle time and under-utilization.
@@ 3.4 Training Details
We employ the GRPO [29] algorithm to train search agents. Specifically, for each input question $x$, $G$ trajectories $\tau_{1},\tau_{2},\cdots,\tau_{G}$ are generated where $\tau_{i}=(s_{0}^{i},a^{i}_{0},s_{1}^{i},\cdots,s_{T_{i}}^{i})$. To optimize the agent, we employ the following loss,
@@
For reward function, we adopt a sparse-reward setting where rewards are computed at trajectory completion. When training from base LLMs, the reward function combines a format reward and F1 score through multiplication. When fine-tuning LRM-based agents (e.g., QwQ), we utilize LLM-as-Judge[20][38] as the reward function and omit format rewards, as these models inherently maintain proper output formatting.
@@ 4.1 Experiment Setup
We first evaluate the agents on single-hop and multi-hop QA tasks. For single-hop questions, we use Natural Questions [15], TriviaQA [12] and PopQA [23]. For multi-hop questions, we use HotpotQA [44], 2WikiMultiHopQA [10], MuSiQue [36], and Bamboogle [28]. We further perform evaluation on more challenging benchmarks including Frames [14], GAIA [24], and xBench-DeepSearch [41] as extra test sets. We evaluate our approach on 1000 randomly sampled instances from the validation sets of HotpotQA, 2WikiMultiHopQA, and MuSiQue. For Bamboogle, Frames, GAIA and xBench-DeepSearch, we use their full test sets. For GAIA, we use the 103 examples from the text-only validation subset [18].
@@
We evaluate the search agents with two settings, each with different types of search tools. In the first setting, local knowledge base with RAG, agents interact with a locally deployed RAG system to retrieve related information from a Wikipedia 2018 corpus [13]. In the other web-based search and browsing setting, agents operate in an interactive web environment with access to both a search engine and a browser tool. For more challenging benchmarks, GAIA, xBench-DeepSearch and Frames, we only conduct evaluations under this web-based setting.
@@
We consider two groups of baselines aligned with the two benchmark categories. For the multi-hop and single-hop QA benchmarks, we include Search-R1(7B/14B/32B) [11], R1-Searcher(7B) [30], Search-o1(QwQ-32B) [18], DeepResearcher [49] and SimpleDeepSearcher [32]. We also prompt Qwen-2.5-7B/32B to directly generate answers without using any tools. On the more challenging benchmarks, we compare against powerful 32B-scale models, including direct generation with QwQ-32B, Search-o1(QwQ-32B) [18], Search-R1-32B [11], WebThinker-QwQ [19],SimpleDeepSearcher-QwQ [32] and WebDancer-32B [39]. All baselines are evaluated using the same tools as our agent to ensure a fair comparison.
@@
We set the turn limit as 32 for 7B and 14B models, and 128 for ASearcher-Web-QwQ. The batch size is set as 128 for 7B and 14B models, and 64 for ASearcher-Web-QwQ. We curate two sets of training data, one for 7B/14B training and the other for QwQ-32B training. These two datasets are both of 35k sizes and open-sourced. Training of ASearcher-Web-QwQ takes approximated 16k H800 GPU hours in total.
Two-Stage Curriculum.
For ASearcher-Web-QwQ, we employ a two-stage curriculum RL training process. In the first stage, the training data covers questions with varying difficulty levels, including simple questions that are solvable with a small number of tool calls. In the second stage, we retain only the questions that require at least 5 tool calls to solve to further activate the long-horizon agentic capability of the model. The final checkpoint of these two stages are name ASearcher-Web-QwQ-v1 and ASearcher-Web-QwQ-v2, respectively.
@@ 4.2 Main Results
As shown in Fig. 8, ASearcher-Web-QwQ obtains +15.0, +22.4, and +14.6 improvements on GAIA, xBench-DeepSearch and Frames respectively. When considering the pass rate, i.e. Pass@4, ASearcher-Web-QwQ also obtains significant gains, especially on xBench-DeepSearch with 24.0 improvements. Significant improvements in pass rate demonstrate that our training pipeline trains the agent to learn complex search strategies to perform precise searches, extract key information, and resolve conflict information.
Table 5: Pass@1 results of ASearcher-Web-QwQ-v2 and baselines, evaluated on GAIA [24], xBench-DeepSearch [41], Frames [14], and HLE-500 [19]. indicates results are obtained from official reports.
Method GAIA xBench-DeepSearch Frames HLE-500
Commercial Deep Research Agents
Kimi-Researcher - 69.0 78.8 26.9
OpenAI DeepResearch 67.0 26.6
General LLMs using Tools
OpenAI-o3 70.5 66.7 84.0 20.2
Qwen3-30B-A3B 35.9 32.0 56.4 13.2
Qwen3-235B-A22B 45.6 46.0 - 20.0
DeepSeek-R1 - 55.0 82.0 24.8
Claude-4-Sonnet 68.3 64.6 80.7 20.3
ASearcher-Web-QwQ (ours)
ASearcher-Web-QwQ-v2 58.7 51.1 74.5 21.5
+ Summary=DeepSeek-V3 60.3 56.4 76.6 23.4
+ Test-time Search (K=16) 71.8 75.0 83.4 24.6
Zero-Shot Transfer with Summary Tool & Test-time Scaling.
Although we adopt a single-model configuration during training time, the agent design of ASearcher is generalizable. Specifically, the webpage summarization process could be supported by external models. We here use DeepSeek-V3 as the webpage summarization model. From Table 5, it is clear that using a more powerful model for summarization improves the performance of ASearcher. We further investigate test-time scaling. Specifically, we run $K=16$ independent runs for each problem in the test set and aggregate the conclusions from these independent runs with DeepSeek-V3. As shown in Table 5, test-time scaling approach leads to competitive performance with commercial systems, including Kimi-Researcher, OpenAI DeepResearch, and OpenAI o3.